# New examples of twisted Brill-Noether loci II

### v1

@@ Abstract.
The authors are members of the research group VBAC (Vector Bundles on Algebraic Curves).
@@ 1. Introduction
Our principal tool in this paper is the dual span construction in its general form (see Subsection 2.1). This allows us to apply results related to Butler’s Conjecture (see Conjecture 1), particularly in the case of line bundles. However, we also consider cases beyond those in which Butler’s Conjecture is known to hold.
@@
In general, for any smooth irreducible projective curve $C$ of genus $g\geq 2$ and $n>1$ the condition that the Brill-Noether number satisfies $\beta(n,d,k)\geq 0$ is neither necessary nor sufficient for the non-emptiness of the Brill–Noether locus $B(n,d,k)$. Nevertheless, for a general curve, we first establish the following result (see Corollary 2.7):
@@
$\text{\it{The same holds for}}\ {B}(n,d,n+1)\ \text{\it{except when }}g=2\ \text{{\it and}}\ d=2n.$
Then, following many partial results (see, for example, [26, 3, 4], and [31]), we prove the following theorem, which completes the proof of Butler’s Conjecture for $(n,d,n+1)$, and also establishes smoothness and irreducibility of the corresponding loci.
@@
We have symmetries $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\cong B^{k}(\mathcal{U}_{2},\mathcal{U}_{1})$, and isomorphisms $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\cong B^{k_{1}}(\mathcal{U}_{1}^{*},\mathcal{U}_{2}^{*}\otimes p_{C}^{*}K_{C})$ and $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\cong B^{k}(\mathcal{U}_{1}\otimes p_{C}^{*}L^{*},\mathcal{U}_{2}\otimes p_{C}^{*}L)$ for any line bundle $L$.
@@
In particular, $B(1,d_{1},k)$ is the classical BN locus $W^{k-1}_{d_{1}}$. If $B(n_{1},d_{1},k)(E_{2})\neq\emptyset$ and $E_{2}$ is stable, then $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\neq\emptyset$. Moreover, if $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\neq\emptyset$ then $\widetilde{B}(n_{1}n_{2},n_{1}d_{2}+n_{2}d_{1},k)\neq\emptyset$.
@@ 2.1. Kernel and dual span bundles
We are particularly interested in cases where $D_{E,V}$ is stable (resp., semistable). With this in mind, we make the following definitions.
(2.3) $S(n,d,v)=\{(E,V)|E\in M(n,d),V\subset H^{0}(E)\text{ generates }E,\\ \dim V=v,D_{E,V}\text{ stable}\};$
(2.4) $\widetilde{S}(n,d,v)=\{(E,V)|[E]\in\widetilde{M}(n,d),V\subset H^{0}(E)\text{ generates }E,\\ \dim V=v,D_{E,V}\text{ semistable}\}.$
@@ 2.2. Rank 11
For $C$ general, considerably more is known. In some cases, we can be more specific and allow $C$ to be an arbitrary Petri curve, that is, a curve for which the morphism $H^{0}(L)\otimes H^{0}(K_{C}\otimes L^{*})\longrightarrow H^{0}(K_{C})$ is injective for every line bundle $L$. For any Petri curve $C$ and any generated line bundle $L$, $D_{L}$ is semistable (in particular, this holds for general $L$ of degree $\geq g+1$) and precise conditions for stability are known [12, Theorem 2]. In fact, this holds for any curve for which $B(1,d,v)=\emptyset$ whenever $\beta(1,d,v)<0$.
@@ 2.3. Non-emptiness of $B(n_{1},d_{1},n_{1}+1)$
Let $C$ be a Petri curve of genus $g\geq 2$. Then $B(n_{1},d_{1},n_{1}+1)\neq\emptyset$ in all the cases indicated in Remark 2.4.
@@ 4.1. The construction
and we want to give conditions for $Im\neq\emptyset$ and $Im\Phi\subset B^{k}(\mathcal{U}_{1},\mathcal{U}_{2}).$

### v2

@@ Abstract.
The authors are members of the research group VBAC (Vector Bundles on Algebraic Curves). The first author thanks Daniel Posada for his help in checking some of the calculations.
@@ 1. Introduction
Our principal tool in this paper is the dual span construction in its general form. This allows us to apply results related to Butler’s Conjecture (see Conjecture 1), particularly in the case of line bundles. However, we also consider cases beyond those in which Butler’s Conjecture is known to hold.
@@
In general, for any smooth irreducible projective curve $C$ of genus $g\geq 2$ and $n>1$ the condition that the Brill-Noether number satisfies $\beta(n,d,k)\geq 0$ is neither necessary nor sufficient for the non-emptiness of the Brill–Noether locus $\widetilde{B}(n,d,k)$. Nevertheless, for a general curve, we first establish the following result (see Corollary 2.7):
@@
[…]
Then, following many partial results (see, for example, [26, 3, 4], and [31]), we prove the following theorem, which completes the proof of Butler’s Conjecture for $(n,d,n+1)$, establishing the smoothness and irreducibility of the corresponding loci.
@@
We have symmetries $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\cong B^{k}(\mathcal{U}_{2},\mathcal{U}_{1})$, and isomorphisms
  • $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\cong B^{k_{1}}(\mathcal{U}_{1}^{*},\mathcal{U}_{2}^{*}\otimes p_{C}^{*}K_{C})$ and
  • $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\cong B^{k}(\mathcal{U}_{1}\otimes p_{C}^{*}L^{*},\mathcal{U}_{2}\otimes p_{C}^{*}L)$ for any line bundle $L$.
@@
In particular, $B(1,d_{1},k)$ is the classical BN locus $W^{k-1}_{d_{1}}$ and if $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\neq\emptyset$ then $\widetilde{B}(n_{1}n_{2},n_{1}d_{2}+n_{2}d_{1},k)\neq\emptyset$. Moreover, $B^{k}(\mathcal{U}_{1},\mathcal{U}_{2})\neq\emptyset$ if $B(n_{1},d_{1},k)(E_{2})\neq\emptyset$ and $E_{2}$ is stable.
@@ 2.1. Kernel and dual span bundles
We are particularly interested in cases where $D_{E,V}$ is stable (resp. semistable). With this in mind, we make the following definitions.
(2.3) $S(n,d,v):=\{(E,V)|E\in M(n,d),V\subset H^{0}(E)\text{ generates }E,\\ \dim V=v,D_{E,V}\text{ stable}\};$
(2.4) $\widetilde{S}(n,d,v):=\{(E,V)|[E]\in\widetilde{M}(n,d),V\subset H^{0}(E)\text{ generates }E,\\ \dim V=v,D_{E,V}\text{ semistable}\}.$
@@ 2.2. Rank 11
For $C$ general, considerably more is known. In some cases, we can be more specific and allow $C$ to be an arbitrary Petri curve, that is, a curve for which the morphism $H^{0}(L)\otimes H^{0}(K_{C}\otimes L^{*})\longrightarrow H^{0}(K_{C})$ is injective for every line bundle $L$. For any Petri curve $C$ and any generated line bundle $L$, $D_{L}$ is semistable (in particular, this holds for general $L$ of degree $\geq g+1$) and precise conditions for stability are known [12, Theorem 2]. In fact, this is still true for any curve for which $B(1,d,v)=\emptyset$ whenever $\beta(1,d,v)<0$.
@@ 2.3. Non-emptiness of $B(n_{1},d_{1},n_{1}+1)$
Let $C$ be a Petri curve of genus $g\geq 2$. Then $B(n_{1},d_{1},n_{1}+1)\neq\emptyset$ in all the cases indicated in Remark 2.4. Note that
$\beta(1,d_{1},n_{1}+1)=\beta(n_{1},d_{1},n_{1}+1)=1$
(resp. $\beta(n_{1},d_{1},n_{1}+1)=0$) if and only if $d_{1}=n_{1}+\frac{n_{1}g}{n_{1}+1}+\frac{1}{n_{1}+1}$ (resp. $d_{1}=n_{1}+\frac{n_{1}}{n_{1}+1})$. For example, for $g=11$, $\beta(2,13,5)=\beta(9,19,10)=1$ and $\beta(10,20,11)=0$. Moreover, $\beta(n_{1},d_{1},n_{1}+a)=1$ if and only if $d_{1}=n_{1}+\frac{an_{1}g}{n_{1}+a}+\frac{a^{2}}{n_{1}+a}$. For example, for $a=2$ and any genus $\beta(2,g+3,4)=1$. Moreover, for a general curve of odd genus $g\geq 5$ by [7, Theorem 5.2(ii), (iii)], $B(2,g+3,4)$ is non-empty.
@@ 3.2. Examples
Note that $\beta(2,9,3)=\beta(8,17,9)=1$ and $\beta(4,12,5)=\beta(9,18,10)=0$.
@@ 4.1. The construction
and we want to give conditions for $Im\Phi\neq\emptyset$ and $Im\Phi\subset B^{k}(\mathcal{U}_{1},\mathcal{U}_{2}).$